OpenStudy (anonymous):

solve 4(3-x)+x=22+2x

OpenStudy (kkutie7):

12-4x+x=22+2x -4x+x-2x=22-12 -3x-2x=10 -5x=10 x=-2

OpenStudy (anonymous):

thank you

OpenStudy (kkutie7):

any time